OpenStudy (anonymous):

2 Math Questions! Need Help!!!!!! Medal!!!! Identify the slope and y-intercept of the graph in the equation. y = 5/4 + 1 A. Slope: 5/4 ; y-intercept: -1 B. Slope: 5/4 ; y-intercept: 1 C. Slope: 4/5 ; y-intercept: -1 D. Slope: 4/5 ; y-intercept: 1 y = -5/4 x + 1 A. Slope: -5/4 ; y-intercept: -1 B. Slope: -5/4 ; y-intercept: 1 C. Slope: -4/5 ; y-intercept: 1 D. Slope: -4/5 ; y-intercept: -1

OpenStudy (anonymous):

@Directrix

OpenStudy (anonymous):

@cwrw238

OpenStudy (anonymous):

ok, well first remember y=mx+b. m is slope, and b is y intercept. does this help?

OpenStudy (anonymous):

can you help me step by step with the first question so I get the hang of it? @DuckDynastyfan923

OpenStudy (anonymous):

ok! so y=5/4x + 1. since 5/4 is where m would be in y=mx+b, this means that 5/4 is your slope. then, since 1 is where b would be in y=mx+b, this means that 1 is your y intercept.

OpenStudy (anonymous):

@Vivid are you still here?

OpenStudy (anonymous):

SO it would be D then?

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

oh wait no it would be B

OpenStudy (anonymous):

Wait,, I meant to put the 1st question as 4/5 not 5/4 @duc

OpenStudy (anonymous):

SO now it would be D correct?

OpenStudy (anonymous):

@DuckDynastyfan923

OpenStudy (anonymous):

no, because it says 4/5. the slope is 5/4.

OpenStudy (anonymous):

The answer is B, because it says that the slope is 5/4, and that the y intercept is 1.

OpenStudy (anonymous):

Wait, The first question is actually: y = 4/5 x + 1 Is it still B? @DuckDynastyfan923

OpenStudy (anonymous):

oh, then it is D.

OpenStudy (anonymous):

ok, cause when i saw it was 4/5 i was sure it wad D. then. So the 2nd Question is A correct? @DuckDynastyfan923

OpenStudy (anonymous):

no, because a shows a negative y intercept. you want one with a positive y intercept, because the one is postitive.

OpenStudy (anonymous):

Ok I get it now thanks @DuckDynastyfan923

OpenStudy (anonymous):

no problem! :)